OpenStudy (anonymous):

Help with high school chemistry 1. This is about calorimetry...Please Help!!

OpenStudy (anonymous):

wts d question

OpenStudy (anonymous):

Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. The water has absorbed the heat of the metal. So, qwater = qmetal

OpenStudy (anonymous):

give the data man

OpenStudy (anonymous):

Part 1: Known Metal: Alluminum Measured mass of metal 27.776 grams Distilled water measurement 25 mL Distilled water temperature 25.2 degrees Celsius Temperature of metal 100.8 degrees Celsius Temperature of mixture 38.9 degrees Celsius

OpenStudy (anonymous):

specific heat capacity of metal???

OpenStudy (anonymous):

The other part says this: Using the formula qmetal = m × c × ΔT, calculate the specific heat of the metal. Use the data from your experiment for the metal in your calculation

OpenStudy (anonymous):

ok wait i'll try it in 2 mins

OpenStudy (anonymous):

u just got2 equate the heat gained by water and heat lost by metal........The calculations are quite tough due to many decimal places......Try it out urself first i hv worked the answer

OpenStudy (anonymous):

Im pretty confused. What should i calculate first?

OpenStudy (anonymous):

its ur wish..........just read the data carefully dts it

OpenStudy (anonymous):

What im asking is, is there an equation I should be following?

OpenStudy (anonymous):

q=energy gained or lost in joules m=27.776 c=specific heat capacity of the sample. [j/(degrees c*grams)] Delta T=the change in temperature when it gains or loses the energy.

OpenStudy (anonymous):

yes just use this and put values in qwater=qmetal

OpenStudy (anonymous):

Im having trouble with [j/(degrees c*grams)] and delta t

OpenStudy (anonymous):

y? take care of units

OpenStudy (anonymous):

Am i using the distilled water temperature or the metal temperature?

OpenStudy (anonymous):

T(initial) for water is distilled one T(final) for water is the mixture temp. T(initial) for metal is metal temp. and T(final) for metal is mixture temp. Gotta go bbye

OpenStudy (anonymous):

27.776*4.18*(25.2-38.9)

OpenStudy (anonymous):

=1,590.620416

OpenStudy (abb0t):

\(q_{water}=mc \Delta T\) start by finding the mass using the known density relationship of water \(Density = \frac{ mass }{ volume }\). Next you can find the relative change in temperature: \(\Delta T = T_{mix}-T_{water}\). Now you have everything to find energy, q. Now, \(q_{water} = - q_{metal}\) so the energy of water is actually negative. Follow the same process for aluminum, get the mass, find the change in temp (∆T). Find, c (specific heat)...and same process for part B.